Find the number of data within 1 standard deviation of the mean: 52, 76, 69, 63, 69, 68, 65, 61.
Thanks!

Answer by jim_thompson5910(35256)

Mean = 523/8 = 65.375

You get the mean by adding up all the numbers and then dividing by 8 (since there are 8 numbers).

-------------------------------------------------------

Standard Deviation = 7.0698

This is approximate. Doing the calculation by hand is tedious so use a calculator here.

-------------------------------------------------------

Lower Bound:
L = (mean) - (standard deviation) = 65.375 - 7.0698 = 58.3052

Upper Bound:
U = (mean) + (standard deviation) = 65.375 + 7.0698 = 72.4448

-------------------------------------------------------

So the data that is roughly within 1 standard deviation of the mean will range from 58.3052 to 72.4448

Those data values that are between 58.3052 and 72.4448 from the data set are: 69, 63, 69, 68, 65, 61

There are 6 data values. So that is the final answer.

If you need more one-on-one help, or if you have any questions, email me at jim_thompson5910@hotmail.com